OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (saifoo.khan):

Try to solve it out!

OpenStudy (anonymous):

how ?..

OpenStudy (saifoo.khan):

Using the formula i used in the last problem.

OpenStudy (anonymous):

how does that work again ? @saifoo.khan sorry to ask.

OpenStudy (saifoo.khan):

\[y-y_1 = m(x-x_1)\]

OpenStudy (saifoo.khan):

Insert the values of m and the point..

OpenStudy (anonymous):

what does that mean ?

OpenStudy (saifoo.khan):

y - (-3) = -6(x-(-2))

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (saifoo.khan):

Do you understand the first step?

OpenStudy (anonymous):

no.

OpenStudy (saifoo.khan):